OpenStudy (anonymous):

Which of the following polynomials has a degree of 3? A. x3y3 - 3xy B. x + y + z C. x2y + 5y2 D. x2y2 - y3

OpenStudy (a1234):

B. Because: x^1 + y^1 + z^1 The sum of those exponents is 3.

OpenStudy (anonymous):

it is c

OpenStudy (a1234):

No, it is not C. C is: x^2y + 5y^2 The sum of those exponents is 4.

OpenStudy (anonymous):

the degree of a polynomial= highest degree of a term of highest degree. A is of 6 degree B is of one degree c is of third degree D is of four degree.

OpenStudy (anonymous):

we have to see the degree of highest degree.

OpenStudy (a1234):

I found this online: If a polynomial has more than one variable, then the degree of that monomial is the sum of the exponents of those variables. In this polynomial, 24xyz, the degree is 3 because the sum of degrees of x, y and z is 1 + 1 + 1 = 3.

OpenStudy (anonymous):

in c, x has degree 2,y of the same term has one degree total=2+1=3

OpenStudy (anonymous):

sum of exponents of one term, not of different terms.

OpenStudy (a1234):

x^2y^1 + 5y^2 The sum is 5.

OpenStudy (anonymous):

24 xyz is of degree 3

OpenStudy (anonymous):

\[x^2y+5y^2,~degree~ of~ highest~ term=2+1=3\]

OpenStudy (anonymous):

\[x^2 y~and~5y^2 ~are~two~different~terms.\]

OpenStudy (a1234):

Ok, you win :).

OpenStudy (anonymous):

it is not a matter of winning or losing ,i just told the reality.

OpenStudy (a1234):

Yes, you told the reality and I told fantasy.

OpenStudy (anonymous):

i don't know where you read all this. we used to teach in this way. according to you ,we are wrong.